# AAAI Press Formatting Instructions for Authors Using — A Guide

### v1

@@ Abstract
Reconstructing objects and extracting high-quality surfaces play a vital role in the real world. Current 4D representations show the ability to render high-quality novel views for dynamic objects but cannot reconstruct high-quality meshes due to their implicit or geometrically inaccurate representations. In this paper, we propose a novel representation that can reconstruct accurate meshes from sparse image input, named Dynamic 2D Gaussians (D-2DGS). We adopt 2D Gaussians for basic geometry representation and use sparse-controlled points to capture 2D Gaussian’s deformation. By extracting the object mask from the rendered high-quality image and masking the rendered depth map, a high-quality dynamic mesh sequence of the object can be extracted. Experiments demonstrate that our D-2DGS is outstanding in reconstructing high-quality meshes from sparse input. More demos and code are available at https://github.com/hustvl/Dynamic-2DGS.
00footnotetext: *Equal contributions.   🖂Corresponding author.
Introduction
@@
The natural world is dynamic, and how to extract the geometric shape of dynamic objects from the sparse input images is an important research problem in computer vision. The main challenges lie in the complex and non-rigid motion, which makes it hard to model accurate motion. Some 3D representations  (Wang et al. 2023b; Huang et al. 2024a) succeed in reconstructing high-quality mesh from static objects but may fall short in 4D objects since we cannot leave every life stock still.
@@
Many approaches succeed in representing dynamic scenes and rendering high-quality novel views. Dynamic NeRFs (Pumarola et al. 2021; Park et al. 2021a, b) mainly suffer from their implicit representations which cause unfriendly memory consumption. Dynamic Gaussian splatting (Wu et al. 2024; Yang et al. 2023a; Huang et al. 2024b) representations can maintain a set of explicit 3D Gaussians (Kerbl et al. 2023) for efficient NVS with high training efficiency, the proposed deformation field shows the potential to modeling object’s motion accurately. However, 3D Gaussians are mainly for NVS and do not have multiview consistency, which causes inaccurate geometry (Huang et al. 2024a). We believe that a proper representation that can reconstruct dynamic objects may achieve two goals: (a) modeling complex motion with high efficiency and (b) enabling the export of high-quality and smooth geometry.
@@
Recently, 2D Gaussian Splatting (2DGS) (Huang et al. 2024a) uses a collection of two-dimensional oriented planar Gaussian disks to represent scenes, ensuring consistent view geometry when modeling surfaces. Extending 2D Gaussians to dynamic object reconstruction and get a high-quality surface is a natural but difficult topic. There are two challenges and the first problem is How to extract a clean consistent surface from sparse input? Geometry floaters also exist in the deformation of 2D Gaussian primitives: the Gaussian owns the same color as the background which is hard to be pruned and would induce inaccurate depth, triggering the low-quality mesh extracted by Truncated Signed Distance Function (TSDF). Moreover, modeling accurate object’s motion also remains a hard issue in the dynamic representations. It’s hard to preserve correspondence between 2D Gaussians’s motion and real-world deformation, which may lead to unstable surfaces. We also observe that while 2D Gaussians benefit from multiview consistency due to their isotropic characteristic, it also degrades Gaussian’s fitting ability when dealing with sparse views.
@@
To tackle the aforementioned problems, we propose to use the sparse-controlled points method (Huang et al. 2024b) to model 2D Gaussians’ motion. Our findings indicate that sparse-controlled points accurately model semi-rigid motion because nearby 2D Gaussian surfels are determined by the sparse-controlled points which provide a more precise canonical-world transformation, ensuring the geometric accuracy of 2D Gaussians in the canonical space and result in smoother surfaces, as illustrated in Fig. LABEL:fig:teaserfig. Furthermore, we introduce a filtering method to remove geometry floaters by filtering the depth image using the rendered RGB mask. This approach allows for an accurate surface mesh to be obtained via the TSDF method.
@@
  • To remove Geometry floaters, we propose to filter the depth image using a mask extracted from the rendered high-quality RGB image.
@@
  • Experiments demonstrate that our D-2DGS achieves state-of-the-art (SOTA) reconstruction quality compared with other advanced representations.
@@
Figure 1: Illustration of Geometry floaters. In the canonical space, all the spatial Gaussian primitives are connected by the control points. After deformation, both objects Gaussians and floater Gaussians are transformed by control points. The similar color to the background makes it hard to be pruned, which causes inaccurate geometry.
Related Works
3D/4D Representations
@@
Representing the 4D world and rendering photo-realistic novel view images has been an important and challenging topic. Explicit representations such as mesh (Collet et al. 2015; Li et al. 2018; Guo et al. 2015; Su et al. 2020), voxels (Guo et al. 2019; Hu et al. 2022; Li et al. 2017), and point clouds enjoy editable manipulations and GPU-friendly applications but suffer from strict multi-view constraints or more sensors. Recently, implicit (Martin-Brualla et al. 2021; Park et al. 2021a, b) representations have demonstrated their efficiency for novel view synthesis from sparse input and many approaches (Xu et al. 2022; Fridovich-Keil et al. 2023; Cao and Johnson 2023; Wang et al. 2023a; Shao et al. 2023; Gan et al. 2023; Li et al. 2022) combine the benefits of these representations. Notably, Gaussian Splatting (Kerbl et al. 2023; Yang et al. 2023b; Wu et al. 2024; Yang et al. 2023a; Zhang et al. 2024) (GS) based representations as an efficient mixture presentation, show the ability to render high-quality novel views and support many downstream tasks (Jiang et al. 2024; Ji et al. 2024; Yi et al. 2024a, b; Yin et al. 2023; Liang et al. 2024). Recent 4D representation, SC-GS (Huang et al. 2024b), proposes a sparse control points method to control the deformation of 3D Gaussians, which reaches outstanding NVS quality and high training efficiency. However, most 3D/4D representations are only for NVS or extract mesh from static scenes and there is less research focus on extracting accurate mesh from dynamic objects in monocular input. Our D-2DGS shows the ability to reconstruct the accurate 3D mesh from 2D images captured in the 4D real world with high efficiency.
Geometry Reconstruction
@@
Geometry reconstructions are mainly from streamable multiview (Collet et al. 2015; Li et al. 2018; Guo et al. 2015; Su et al. 2020) constraints. Recently, 3DGS (Kerbl et al. 2023) is proven to enjoy high training efficiency and shows great potential to reconstruct high-quality mesh from spatial Gaussians. Many works are proposed to improve the geometric accuracy of 3DGS. SuGaR (Guédon and Lepetit 2024) introduces a regularization term that encourages the Gaussians to align well with the surface of the scene. NeuSG (Chen, Li, and Lee 2023) jointly optimizes 3DGS and neural implicit models, combining the advantages of both, enabling the generation of complete surfaces with high details. Similarly, 3DGSR (Lyu et al. 2024) also encourages the alignment of 3D Gaussian and SDF to improve the surface reconstruction effect. 2D Gaussian Splatting (2DGS) (Huang et al. 2024a)is a breakthrough in mesh reconstruction since the view geometry consistency is maintained by planar Gaussian disks.
@@
The above geometric reconstruction methods are all for static scenes. Recently, some works have been attempted on the geometric reconstruction of dynamic scenes. Dynamic Gaussian Mesh (DG-Mesh) (Liu, Su, and Wang 2024) aims to reconstruct a high-fidelity and temporally consistent mesh from a given monocular video. DG-Mesh introduces Gaussian grid anchoring to encourage uniformly distributed Gaussian distributions and obtains better grid reconstruction by grid densification and pruning of deformed Gaussian distributions. MaGS (Ma, Luo, and Yang 2024) propose mutual-adsorbed mesh Gaussian representation to refine the generated mesh. Vidu4D (Wang et al. 2024b) uses Dynamic Gaussian Surfels as scene representation primitives, achieving text to 4D generation with high-quality appearance and geometry. Our methods propose to model Gaussian surfels’ motion by sparse-controlled points, maintaining consistent and smooth surface mesh generation.
Preliminaries
2D Gaussian Splatting
@@
$\mathcal{G}(\mathbf{u})=exp(-\frac{u^{2}+v^{2}}{2}).$ (2)
@@
$c(x)=\sum_{i=1}\mathbf{c}_{i}\alpha\mathcal{G}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-\alpha_{j}\mathcal{G}(\mathbf{u}(x))).$ (3)
However, 2D Gaussians can only reconstruct 3D mesh from static scenes. We propose to use Sparse-Controlled points to model 2D Gaussians deformation and extend 2D Gaussians representing the 4D world successfully.
Sparse-Controlled Gaussian Splatting
Sparse-Controlled Gaussian Splatting (SC-GS) (Huang et al. 2024b) shows excellent performance in the task of synthesizing novel views in dynamic scenes. It uses sparse control points to control dense 3D Gaussians in the scene. For each control point at timestep $t$, SC-GS uses MLP $\Phi$ to predict a rotation matrix $R_{i}^{t}\in\mathbf{SE}(3)$ and rotation matrix $T_{i}^{t}\in\mathbb{R}^{3}$. For each Gaussian $G_{j}:(\mu_{j},q_{j},s_{j},\sigma_{j},sh_{j})$, use k-nearest neighbor (KNN) search to get the $K$ neighboring control points $\{p_{i},i=0,1,...,k\}$. SC-GS employs LBS (Sumner, Schmid, and Pauly 2007) $\mathcal{I}$ to compute the warped Gaussian $\mu_{j}^{t}$ and $q_{j}^{t}$:
$\mu_{j}^{t}=\sum_{k\in\mathcal{N}_{j}}w_{jk}(R_{k}^{t}(\mu_{j}-p_{k}))+p_{k}+T_{k}^{t}),$ (4)
$q_{j}^{t}=(\sum_{k\in\mathcal{N}_{j}}w_{jk}r_{k}^{t})\otimes q_{j},$ (5)
@@
$w_{jk}=\frac{\hat{w}_{jk}}{\sum_{k\in\mathcal{N}_{j}}\hat{w}_{jk}},\mathrm{where}~{}\hat{w}_{jk}=exp(-\frac{d_{jk}^{2}}{2o_{k}^{2}}),$ (6)
@@
However, SC-GS (Huang et al. 2024b) mainly chooses 3D-GS (Kerbl et al. 2023) as its geometry representation, leading to multiview inconsistency. Meanwhile, There are fewer geometry constraints during the optimization of canonical Gaussians, we mainly choose 2DGS (Huang et al. 2024a) as basic geometry, and propose an outlier removal method to overcome the Geometry floaters as shown in Fig. 1 from the sparse monocular input.
Method
Figure 2: Framework of our D-2DGS. Sparse points are bonded with canonical 2D Gaussians. Deformation networks are used to predict each sparse control point’s control signals given any timestamp. The image and depth are rendered by deformed 2D Gaussians with alpha blending. To get high-quality meshes, depth images are filtered by rendered images with RGB mask, and then TSDF is applied on multiview depth images and RGB images.
Dynamic 2D Gaussians Framework
The pipeline of our framework is illustrated in Fig. 2. We adopt 2D Gaussian primitives $\mathcal{G}$ for scene representation. We follow deformation-based Gaussian splatting to compute the deformation of each 2D Gaussians at timestamp $t_{j}$:
$\mathcal{G}^{j}=Deform(\mathcal{G},t_{j},\mathbf{P};\Phi),$ (7)
where $Deform$ denotes deformation process, $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ denotes sparse control points. Then $\mathcal{G}^{j}$ are used to render the RGB image $\mathbf{I}$ and the depth image $\mathbf{D}$ by differential splatting algorithm (Huang et al. 2024b; Yifan et al. 2019) $\mathcal{S}$ with a camera parameters $[\mathbf{R},\mathbf{t}]$:
$\mathbf{I},\mathbf{D}=\mathcal{S}\left(\mathcal{G}^{j};\theta),[\mathbf{R},\mathbf{t}]\right).$ (8)
@@
$\mathbf{D}^{\prime}=Flitering(\mathbf{I},\mathbf{D}),$ (9)
@@
Sparse point controlled 2D Gaussians

### v2

# Dynamic 2D Gaussians: Geometrically Accurate Radiance Fields for Dynamic Objects

@@ Abstract.
Reconstructing objects and extracting high-quality surfaces play a vital role in the real world. Current 4D representations show the ability to render high-quality novel views for dynamic objects, but cannot reconstruct high-quality meshes due to their implicit or geometrically inaccurate representations. In this paper, we propose a novel representation that can reconstruct accurate meshes from sparse image input, named Dynamic 2D Gaussians (D-2DGS). We adopt 2D Gaussians for basic geometry representation and use sparse-controlled points to capture the 2D Gaussian’s deformation. By extracting the object mask from the rendered high-quality image and masking the rendered depth map, we remove floaters that are prone to occur during reconstruction and can extract high-quality dynamic mesh sequences of dynamic objects. Experiments demonstrate that our D-2DGS is outstanding in reconstructing detailed and smooth high-quality meshes from sparse inputs. The code is available at https://github.com/hustvl/Dynamic-2DGS.
Gaussian Splatting; 4D Reconstruction; Sparse Perspective
submissionid: 3241journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755328isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer graphics
Figure 1. Applications of D-2DGS. Given sparse 2D dynamic images, D-2DGS can reconstruct dynamic objects and extract high-quality normal, time-aware mesh, render quality depth, and novel view images.
1. Introduction
@@
The natural world is dynamic, and how to extract the geometric shape of dynamic objects from the sparse input images is an important research problem in computer vision. The main challenges lie in the complex and non-rigid motion, which makes it hard to model motion accurately. Some 3D representations  (Wang et al., 2023; Huang et al., 2024b) succeed in reconstructing high-quality mesh from static objects but may fall short in 4D objects since we cannot leave every life stock still.
@@
Many approaches succeed in representing dynamic scenes and rendering high-quality novel views. Dynamic NeRFs (Pumarola et al., 2021; Park et al., 2021a, b) mainly suffer from their implicit representations, which cause unfriendly memory consumption. Dynamic Gaussian splatting (Wu et al., 2024; Yang et al., 2024a; Huang et al., 2024a) representations can maintain a set of explicit 3D Gaussians (Kerbl et al., 2023) for efficient NVS with high training efficiency, and the proposed deformation field shows the potential to model an object’s motion accurately. However, 3D Gaussians are mainly for NVS and do not have multiview consistency, which causes inaccurate geometry (Huang et al., 2024b). Some 3D Gaussian-based approaches also succeed in reconstructing a 3D mesh from dynamic images. We also found it (Liu et al., 2025) lacks high-quality details. We believe that a proper representation that can reconstruct dynamic objects may achieve two goals: (a) modeling complex motion with high efficiency and (b) enabling the export of high-quality and smooth geometry.
@@
Recently, 2D Gaussian Splatting (2DGS) (Huang et al., 2024b) uses a collection of two-dimensional oriented planar Gaussian disks to represent scenes, ensuring consistent view geometry when modeling surfaces. Extending 2D Gaussians to dynamic object reconstruction and get a high-quality surface is a natural but difficult topic. There are two challenges, and the first problem is How to extract a clean, consistent surface from sparse input? Geometry floaters also exist in the deformation of 2D Gaussian primitives: the Gaussian owns the same color as the background, which is hard to be pruned and would induce inaccurate depth, triggering the low-quality mesh extracted by Truncated Signed Distance Function (TSDF). Moreover, modeling an accurate object’s motion also remains a hard issue in the dynamic representations. It’s hard to preserve correspondence between 2D Gaussians’ motion and real-world deformation, which may lead to unstable surfaces. Meanwhile, We observe that while 2D Gaussians benefit from multiview consistency due to their isotropic characteristic, it also degrades the Gaussian’s fitting ability when dealing with sparse views.
@@
To tackle the aforementioned problems, we propose to use the sparse-controlled points method (Huang et al., 2024a) to model 2D Gaussians’ motion. Our findings indicate that sparse-controlled points accurately model semi-rigid motion because nearby 2D Gaussian surfels are determined by the sparse-controlled points, which provide a more precise canonical-world transformation, ensuring the geometric accuracy of 2D Gaussians in the canonical space and resulting in smoother surfaces. Furthermore, we introduce a filtering method to remove geometry floaters by filtering the depth image using the rendered high-quality RGB mask. This approach allows for an accurate surface mesh to be obtained via the TSDF method.
@@
  • To extract a clean mesh, it is necessary to remove the Geometry floaters generated during reconstruction. Therefore, we propose to filter the depth image using a mask extracted from the rendered high-quality RGB image.
@@
  • Experiments demonstrate that our D-2DGS achieves state-of-the-art (SOTA) reconstruction/mesh rendering quality compared with other advanced representations.
@@
Figure 2. Illustration of Geometry floaters. After deformation, both objects Gaussians and floater Gaussians are transformed by control points. The similar color to the background makes it hard to be pruned, which causes inaccurate geometry.
2. Related Works
2.1. 3D/4D Representations
@@ 2. Related Works
Representing the 4D world and rendering photo-realistic novel view images has been an important and challenging topic. Explicit representations such as mesh (Li et al., 2018; Guo et al., 2015), voxels (Guo et al., 2019; Hu et al., 2022; Li et al., 2017), and point clouds (Xu et al., 2022) enjoy editable manipulations and GPU-friendly applications but suffer from strict multi-view constraints or more sensors. Recently, implicit (Martin-Brualla et al., 2021; Park et al., 2021a, b) representations have demonstrated their efficiency for novel view synthesis from sparse input, and many approaches (Fridovich-Keil et al., 2023; Cao and Johnson, 2023; Shao et al., 2023; Fang et al., 2022) combine the benefits of these representations. Notably, Gaussian Splatting (Kerbl et al., 2023; Yang et al., 2024b; Wu et al., 2024; Yang et al., 2024a; Zhang et al., 2024; Lei et al., 2025) (GS) based representations as an efficient mixture presentation, show the ability to render high-quality novel views and support many downstream tasks (Jiang et al., 2024; Ji et al., 2024; Yi et al., 2024a, b). Recent 4D representation, SC-GS (Huang et al., 2024a), proposes a sparse control points method to control the deformation of 3D Gaussians, which reaches outstanding NVS quality and high training efficiency. However, most 3D/4D representations are only for NVS or extracting mesh from static scenes, and there is less research focus on extracting an accurate mesh from dynamic objects. Our D-2DGS shows the ability to reconstruct an accurate 3D mesh from 2D images captured in the 4D world.
2.2. Geometry Reconstruction
@@
Geometry reconstructions are mainly from streamable multiview constraints (Collet et al., 2015; Li et al., 2018; Guo et al., 2015). Recently, 3DGS (Kerbl et al., 2023) is proven to enjoy high training efficiency and shows great potential to reconstruct high-quality mesh from spatial Gaussians. Many works are proposed to improve the geometric accuracy of 3DGS. SuGaR (Guédon and Lepetit, 2024) introduces a regularization term that encourages the Gaussians to align well with the surface of the scene. NeuSG (Chen et al., 2023) jointly optimizes 3DGS and neural implicit models, combining the advantages of both, enabling the generation of complete surfaces with high details. Similarly, 3DGSR (Lyu et al., 2024) also encourages the alignment of 3D Gaussian and SDF to improve the surface reconstruction effect. 2D Gaussian Splatting (2DGS) (Huang et al., 2024b)is a breakthrough in mesh reconstruction since the view geometry consistency is maintained by planar Gaussian disks. Recently, 3D foundation models  (Xiang et al., 2025; Tochilkin et al., 2024; Li et al., 2025; Ye et al., 2025) enable high-quality mesh generation from sparse input.
@@
The above geometric reconstruction methods are all for static scenes. Recently, some works have been attempted on the geometric reconstruction of dynamic scenes. Dynamic Gaussian Mesh (DGMesh) (Liu et al., 2025) aims to reconstruct a high-fidelity and temporally consistent mesh from a given monocular video. DGMesh introduces Gaussian grid anchoring to encourage uniformly distributed Gaussian distributions and obtains better grid reconstruction by grid densification and pruning of deformed Gaussian distributions. MaGS (Ma et al., 2024) proposes a mutual-adsorbed mesh Gaussian representation to refine the generated mesh. Vidu4D (Wang et al., 2024b) uses Dynamic Gaussian Surfels as scene representation primitives, achieving text-to-4D generation with high-quality appearance and geometry. Here are many concurrent works (Ma et al., 2024; Li et al., 2024) that propose to use a 3D Gaussian representation to reconstruct a mesh from dynamic scenes. Papers(Cai et al., 2024; Wang et al., 2024a) adopt a deformation field network and Gaussian surfels to model the scene’s deformation. Our methods propose to model Gaussian surfels’ motion by sparse-controlled points, maintaining consistent and smooth surface mesh generation.
3. Preliminaries
3.1. 2D Gaussian Splatting
@@ 3. Preliminaries
(2) $\mathcal{G}(\mathbf{u})=exp(-\frac{u^{2}+v^{2}}{2}).$
@@
(3) $c(x)=\sum_{i=1}\mathbf{c}_{i}\alpha\mathcal{G}(\mathbf{u}(x))\prod_{j=1}^{i-1}(1-\alpha_{j}\mathcal{G}(\mathbf{u}(x))).$
However, 2D Gaussians can only reconstruct a 3D mesh from static scenes. We propose to use Sparse-Controlled points to model 2D Gaussians deformation and extend 2D Gaussians to represent the 4D world successfully.
Figure 3. Framework of our D-2DGS. Sparse points are bonded with canonical 2D Gaussians. Deformation networks are used to predict each sparse control point’s control signals given any timestamp. The image and depth are rendered by deformed 2D Gaussians with alpha blending. To get high-quality meshes, depth images are filtered by rendered images with RGB mask, and then TSDF is applied on multiview depth images and RGB images.
3.2. Sparse-Controlled Gaussian Splatting
Sparse-Controlled Gaussian Splatting (SC-GS) (Huang et al., 2024a) shows excellent performance in the task of synthesizing novel views in dynamic scenes. It uses sparse control points to control dense 3D Gaussians in the scene. For each control point at timestep $t$, SC-GS uses MLP $\Phi$ to predict a rotation matrix $R_{i}^{t}\in\mathbf{SE}(3)$ and translation matrix $T_{i}^{t}\in\mathbb{R}^{3}$. For each Gaussian $G_{j}:(\mu_{j},q_{j},s_{j},\sigma_{j},sh_{j})$, use k-nearest neighbor (KNN) search to get the $K$ neighboring control points $\{p_{i},i=0,1,...,k\}$. SC-GS employs LBS (Sumner et al., 2007) $\mathcal{I}$ to compute the warped Gaussian $\mu_{j}^{t}$ and $q_{j}^{t}$:
(4) $\mu_{j}^{t}=\sum_{k\in\mathcal{N}_{j}}w_{jk}(R_{k}^{t}(\mu_{j}-p_{k}))+p_{k}+T_{k}^{t}),$
(5) $q_{j}^{t}=(\sum_{k\in\mathcal{N}_{j}}w_{jk}r_{k}^{t})\otimes q_{j},$
@@ 3.2. Sparse-Controlled Gaussian Splatting
(6) $w_{jk}=\frac{\hat{w}_{jk}}{\sum_{k\in\mathcal{N}_{j}}\hat{w}_{jk}},\mathrm{where}~\hat{w}_{jk}=exp(-\frac{d_{jk}^{2}}{2o_{k}^{2}}),$
@@
However, SC-GS (Huang et al., 2024a) mainly chooses 3D-GS (Kerbl et al., 2023) as its geometry representation, leading to multiview inconsistency. Meanwhile, there are fewer geometry constraints during the optimization of canonical Gaussians, we mainly choose 2DGS (Huang et al., 2024b) as basic geometry, and propose an outlier removal method to overcome the Geometry floaters as shown in Fig. 2.
4. Method
4.1. Dynamic 2D Gaussians Framework
The pipeline of our framework is illustrated in Fig. 3. We adopt 2D Gaussian primitives $\mathcal{G}$ for scene representation. We follow deformation-based Gaussian splatting to compute the deformation of each 2D Gaussian at timestamp $t_{j}$:
(7) $\mathcal{G}^{j}=\Phi(\mathcal{G},t_{j},\mathbf{P}),$
where $\Phi$ denotes deformation network, $\mathbf{P}=\{p_{i}\}_{i=1}^{N}$ denotes sparse control points. Then $\mathcal{G}^{j}$ are used to render the RGB image $\mathbf{I}$ and the depth image $\mathbf{D}$ by differential splatting algorithm (Huang et al., 2024a; Yifan et al., 2019) $\mathcal{S}$ with a camera parameters $\mathbf{R},\mathbf{t}]$:
(8) $\mathbf{I},\mathbf{D}=\mathcal{S}\left((\mathcal{G}^{j};\theta),[\mathbf{R},\mathbf{t}]\right).$
@@ 4.1. Dynamic 2D Gaussians Framework
(9) $\mathbf{D}^{\prime}=Flitering(\mathbf{I},\mathbf{D}),$
@@
4.2. Sparse point controlled 2D Gaussians